OpenStudy (anonymous):

Is sqrt(1-sin^2Θ)=cosΘ true?

OpenStudy (anonymous):

\[\huge \sqrt{(1-\sin^2Θ)} =\sqrt{\cos^2Θ} = \cos \theta\] thus it is true.

OpenStudy (anonymous):

Ok thank you. Can you help me figure out which quadrant it lies in?

OpenStudy (anonymous):

if you provide some boundaries then i may help you..

OpenStudy (anonymous):

This is all I have to go off of. Since cosine is positive, I think it is Quadrant II or Quadrant III. Am I correct?

OpenStudy (anonymous):

@ganeshie8 can you tell me if I'm right please?

ganeshie8 (ganeshie8):

hint : Cosine is positive in I and IV quadrants

ganeshie8 (ganeshie8):

So, it must lie in I and IV quadrants right ?

OpenStudy (anonymous):

Oh ok. Is cosine x or y?

ganeshie8 (ganeshie8):

cosine is x

ganeshie8 (ganeshie8):

sine is y

OpenStudy (anonymous):

Ok I see where I went wrong now. Thank you!

ganeshie8 (ganeshie8):

good :) u wlc !

OpenStudy (anonymous):

it is not true if cosine is negative

OpenStudy (anonymous):

\[\large \sqrt{1-\sin^2(\frac{2\pi}{3})}\neq \cos(\frac{2\pi}{3})\] for example

OpenStudy (anonymous):

I'm confused now. So that means sqrt(1-sin^2Θ)=cosΘ is false then?

OpenStudy (anonymous):

\[\sqrt{1-\sin^2(x)}\] is always positive

OpenStudy (anonymous):

but cosine is not always positive

ganeshie8 (ganeshie8):

it is true always cuz sqrt is always postive

ganeshie8 (ganeshie8):

however the converse is not true always, but that shouldnt matter here

OpenStudy (anonymous):

what is true is that \[\cos(\theta)=\pm\sqrt{1-\sin(\theta)}\]

ganeshie8 (ganeshie8):

also, if u live in I and IV quadrants, you dont have to bother cuz cos(x) is always positive

OpenStudy (anonymous):

Ok because sqrt(1-sin^2Θ)=sqrt(sin^2Θ+cos2Θ-sin2Θ=cosΘ) right?

OpenStudy (anonymous):

Ok wait, I get what @satellite73 is saying now

ganeshie8 (ganeshie8):

that looks bit convincing, but satellite is telling an important concept... oh good :)

OpenStudy (anonymous):

Thank you both :). You explained it a lot better than my teacher

ganeshie8 (ganeshie8):

glad to hear :)